# Structured Agent Distillation for Large Language Model

### v1

@@ 1 Introduction
Despite their effectiveness, LLM-based general agents remain costly to deploy due to model size and inference overhead. To address this, recent work distills large agents into smaller student models. However, most approaches rely on token-level supervision hinton2015distilling ; turc2019well ; eldan2023tinystories ; gu2024minillm ; minixhofer2025cross ; cui2025multi ; zhang2024knowledge , which treats the agent trajectory as a flat token sequence and aligns predictions step by step—ignoring its structured composition of reasoning and action.
@@ 7 Discussion
Future Work. We aim to extend span-aware supervision to multimodel agents, and explore alignment for more complex structures, such as hierarchical subgoals or latent plans. We plan to use diffusion models meng2024instructgie to generate small samples liu2023interpretable ; liu2021explainable ; liu2022efficient for multimodal agent research in healthcare liu2024brain ; chinta2023optimization ; wang2025graph ; chinta2024ai . These samples will be combined with large language models (LLMs) for downstream tasks such as 3D reconstruction lei2023mac ; dong2024physical ; dong2024df . In parallel, we will benchmark recent compression techniques tan2025perturbation ; liu2024tsla ; liu2025rora ; liu2024toward ; limutual ; tan2025harmony ; ji2025computation ; liu2025toward to support efficient deployment on embedded devices yuan2021work ; yuan2022mobile ; liu2024tsla ; ji2025computation ; liu2023scalable .
@@ Appendix C Evaluation Metrics
    CoT Match $\text{CoT Match Rate}=\frac{1}{N}\sum_{i=1}^{N}\frac{|\tau_{S}^{(r,i)}\cap\tau_{T}^{(r,i)}|}{|\tau_{T}^{(r,i)}|},$
@@
    CoT Match $\text{CoT Match Rate}=\frac{6}{9}=66.67\%.$
@@
    Average $\text{Average Steps}=\frac{1}{N}\sum_{i=1}^{N}\text{Steps}(\tau^{(i)}),$ (11)
@@
    Average $\text{Average Steps}=\frac{16+13+20}{3}=16.33.$
@@ D.3 Example and Loss Application
[…]

### v2

@@ 1 Introduction
Despite their effectiveness, LLM-based general agents remain costly to deploy due to model size and inference overhead. To address this, recent work distills large agents into smaller student models. However, most approaches rely on token-level supervision hinton2015distilling ; liu2025toward ; liu2025rora ; li2025mutual ; liu2025rcr ; yang2025fairsmoe ; eldan2023tinystories ; gu2024minillm ; minixhofer2025cross ; cui2025multi ; zhang2024knowledge , which treats the agent trajectory as a flat token sequence and aligns predictions step by step—ignoring its structured composition of reasoning and action.
@@ 7 Discussion
Future Work. We aim to extend span-aware supervision to multimodel agents, and explore alignment for more complex structures, such as hierarchical subgoals or latent plans. We plan to use diffusion models meng2024instructgie to generate small samples liu2023interpretable ; liu2021explainable ; liu2022efficient for multimodal agent research in healthcare chinta2025ai ; liu2024brain ; chinta2023optimization ; wang2025graph . These samples will be combined with large language models (LLMs) for downstream tasks such as 3D reconstruction lei2023mac ; dong2024physical ; dong2024df . In parallel, we will benchmark recent compression techniques tan2025perturbation ; liu2025rora ; li2025mutual ; tan2025harmony ; ji2025computation ; liu2025toward to support efficient deployment on embedded devices niu2025mobile ; yuan2021work ; yuan2022mobile ; ji2025computation ; liu2023scalable .
@@ Appendix C Evaluation Metrics
    $\text{CoT Match Rate}=\frac{1}{N}\sum_{i=1}^{N}\frac{|\tau_{S}^{(r,i)}\cap\tau_{T}^{(r,i)}|}{|\tau_{T}^{(r,i)}|},$
@@
    $\text{CoT Match Rate}=\frac{6}{9}=66.67\%.$
@@
    $\text{Average Steps}=\frac{1}{N}\sum_{i=1}^{N}\text{Steps}(\tau^{(i)}),$ (11)
@@
    $\text{Average Steps}=\frac{16+13+20}{3}=16.33.$
@@ D.3 Example and Loss Application
$x=[\texttt{Reasoning},:,\texttt{I},\texttt{should},\texttt{find},\texttt{the},\texttt{fridge},.,\texttt{It},\texttt{'s},\texttt{likely},\texttt{in},\texttt{the},\texttt{kitchen},.,\texttt{Action},:,\texttt{goto},\texttt{kitchen}]$